# thanks: These authors contributed equallythanks: These authors contributed equally

### v2

@@ I Introduction
Figure 2: Interaction energy ($\varDelta E$) in five lattice polymer models. The range of interactions is denoted by $d$; for $d=0$ or $1$, the interaction is between two segments occupying the same site or nearest-neighbor sites, respectively. The RW has no interactions between segments. The SAW introduces a hard-core repulsion ($d=0$) that prohibits segment overlap. The NAW introduces an additional hard-shell repulsion ($d=1$) between nearest-neighbor segments. The DJ model and the ISAW incorporate soft-core ($d=0$) and soft-shell ($d=1$) repulsions, respectively, with interaction strength $\varepsilon\,(>0)$. Varying $\varepsilon$ in these models allows continuous crossovers between the RW, SAW, and NAW.
@@ II Models: RW–SAW and SAW–NAW crossovers
where $\varepsilon\,(>0)$ is the repulsive interaction energy (bottom panels of Fig. 2), $m=m(\omega)$ is the number of interacting segment pairs, and $\omega$ is the configuration of the RW (SAW) for the DJ model (ISAW). In the DJ model, $m$ increases whenever multiple segments occupy the same site. When $v\,(\geq 2)$ segments overlap at one site, $m$ increases by the binomial coefficient $\binom{v}{2}\equiv v(v-1)/2$. In the ISAW, an interacting segment pair is defined as two sequentially nonadjacent segments occupying nearest-neighbor sites.
@@
$S(r,\beta\varepsilon)$ $=$ $k_{B}\beta\left[A(r,\beta\varepsilon)-U(r,\beta\varepsilon)\right],$ (5)
@@
$k(r,\beta\varepsilon)$ $\equiv$ $\frac{1}{\beta}\left[\left(\frac{1}{Z(r,\beta\varepsilon)}\sum_{m=0}^{m_{\mathrm{ub}}}\frac{\varDelta W_{n,m}(r)}{\varDelta r}\,e^{-\beta\varepsilon m}\right)^{2}\right$. (6)
$\left.\quad-\frac{1}{Z(r,\beta\varepsilon)}\sum_{m=0}^{m_{\mathrm{ub}}}\frac{\varDelta^{2}W_{n,m}(r)}{\varDelta r^{2}}\,e^{-\beta\varepsilon m}\right],$
where $\varDelta W_{n,m}(r)\equiv\left[W_{n,m}(r+\varDelta r)-W_{n,m}(r-\varDelta r)\right]/2$ and $\varDelta^{2}W_{n,m}(r)\equiv W_{n,m}(r+\varDelta r)-2W_{n,m}(r)+W_{n,m}(r-\varDelta r)$. Here, $\varDelta r\equiv 2$ because $\omega$ exists only when $r$ and $n$ are both even or both odd. The stiffness $k$ is the sum of energetic ($k_{U}$) and entropic ($k_{S}$) contributions, where $k_{U}(r,\beta\varepsilon)\equiv\partial^{2}U(r,\beta\varepsilon)/\partial r^{2}$ and $k_{S}(r,\beta\varepsilon)\equiv-T\partial^{2}S(r,\beta\varepsilon)/\partial r^{2}$. We use
@@ 99 $2069333612$ $32887440$ $204330$
$W_{n}^{\mathrm{RW}}(r)=\binom{n}{\frac{n-r}{2}}\sum_{k=0}^{\frac{n-r}{2}}\binom{\frac{n-r}{2}}{k}\binom{\frac{n+r}{2}}{k}\binom{2k}{k}$. (8)
@@ V Emergence of negative energetic elasticities in lattice polymer chains
Figure 3 shows $\beta k$, $\beta k_{S}$, $\beta k_{U}$, $k_{U}/k$, and $\tau\beta\varepsilon$ for the DJ model and ISAW for $(n,r)=(20,10)$. The left column of Fig. 3 presents these values for the DJ model, calculated using Table 1, which enables the computation of the differences $\varDelta W_{n,m}(r)$ and $\varDelta^{2}W_{n,m}(r)$ at $r=10$ in Eq. (6). The two crossovers interconnecting the RW, SAW, and NAW ($\beta k_{U}=0$) reveal that both the DJ model and ISAW exhibit negative energetic elasticity ($\beta k_{U}<0$) with $|k_{U}/k|$ maximized around $\beta\varepsilon\sim 1$. For $(n,r)=(20,10)$ and the same $\beta\varepsilon$, $k^{\mathrm{DJ}}$ is $4.6$ to $8.7$ times larger than $k^{\mathrm{ISAW}}$ because the configuration space of the RW is larger than that of the SAW. The ratio of $|(k_{U}/k)^{\mathrm{DJ}}|$ to $|(k_{U}/k)^{\mathrm{ISAW}}|$ ranges from $1.1$ to $2.7$, demonstrating the enhanced negative energetic elasticity in the DJ model compared to that in the ISAW.
@@ Appendix A Upper bounds on the number of interacting segment pairs
For the DJ model, the tightest upper bound on (i.e., the maximum value of) $m$ for a given chain length $n$ and end-to-end distance $r$ is achieved when the overlaps of segments are restricted to the minimum number of sites. This is because the number of interacting segment pairs at a single site increases quadratically with the number of overlapping segments $v$ at that site, according to the binomial coefficient $\binom{v}{2}\equiv v(v-1)/2$ for $v\geq 2$. To construct a configuration that maximizes $m$ for a given $n$ and $r$, we arrange the initial $r$ segments to reach the site $(r,0,0)$, and then arrange the remaining $s\equiv n-r$ segments to oscillate back and forth between two sites along the $x$-axis. For $r=0$, the oscillation is between the sites $(1,0,0)$ and $(0,0,0)$, and for $r\geq 1$, the oscillation is between (r1,0,0)(r-1,0,0) and $(r,0,0)$, yielding the maximum of $m$ for given $n$ and $r$:
$m_{\mathrm{ub}}^{\mathrm{RW}}=\begin{cases}\displaystyle\left\lfloor\frac{n^{2}}{4}\right\rfloor=\left\lfloor\frac{s^{2}}{4}\right\rfloor&(r=0),\\ \\ \displaystyle\left\lfloor\frac{(n-r+1)^{2}}{4}\right\rfloor=\left\lfloor\frac{(s+1)^{2}}{4}\right\rfloor&(r\geq 1).\end{cases}$ (11)
Here, $\lfloor x\rfloor$ is the floor function, which represents the largest integer less than or equal to $x$. Equation (11) gives the maximum of $m$ that satisfies $W^{\mathrm{RW}}_{n,m}(n-s)\geq 1$ in Tables S1 to S18 in the Supplemental Material [27]. In the summation of Eq. (6) for the DJ model, $m_{\mathrm{ub}}^{\mathrm{RW}}$ depends on $r$ for a fixed $n$. Thus, the maximum of $m_{\mathrm{ub}}^{\mathrm{RW}}$, corresponding to the $r-\varDelta r$ case, should be used.
@@ Appendix B Polynomials for the number of random walks with small slack
$W^{\mathrm{RW}}_{n,0}(n-8)$ $=$ $\frac{1}{288}(35n^{8}-1620n^{7}+33606n^{6}-407160n^{5}+3151827n^{4}-15991140n^{3}+52024708n^{2}$ (41)
@@
$W^{\mathrm{RW}}_{n,1}(n-8)$ $=$ $\frac{2}{9}(10n^{7}-383n^{6}+6235n^{5}-54875n^{4}+270421n^{3}-654914n^{2}+234858n$ (42)
@@
$W^{\mathrm{RW}}_{n,2}(n-8)$ $=$ $\frac{1}{18}(10n^{7}-182n^{6}+466n^{5}+5671n^{4}+53494n^{3}-1346021n^{2}+7741074n$ (43)
@@
$W^{\mathrm{RW}}_{n,0}(n-10)$ $=$ $\frac{1}{3600}(63n^{10}-4585n^{9}+153060n^{8}-3081950n^{7}+41448599n^{6}-389383445n^{5}+2591570990n^{4}$ (63)
@@
$W^{\mathrm{RW}}_{n,1}(n-10)$ $=$ $\frac{1}{144}(70n^{9}-4395n^{8}+122592n^{7}-1976910n^{6}+20015526n^{5}-127970403n^{4}+476189444n^{3}$ (64)
@@
$W^{\mathrm{RW}}_{n,2}(n-10)$ $=$ $\frac{1}{288}(35n^{9}-970n^{8}-1738n^{7}+356852n^{6}-4644373n^{5}+7548302n^{4}+338801436n^{3}$ (65)
@@
$W^{\mathrm{RW}}_{n,3}(n-10)$ $=$ $\frac{2}{9}(10n^{8}-481n^{7}+10542n^{6}-145552n^{5}+1475385n^{4}-11407483n^{3}+62293463n^{2}$ (66)
@@
$W^{\mathrm{RW}}_{n,4}(n-10)$ $=$ $\frac{1}{18}(5n^{8}+112n^{7}-11966n^{6}+304687n^{5}-4138721n^{4}+34852933n^{3}-185222134n^{2}$ (67)
@@
$W^{\mathrm{RW}}_{n,5}(n-10)$ $=$ $\frac{1}{90}(370n^{7}-7025n^{6}-90077n^{5}+3908575n^{4}-47417645n^{3}+270658210n^{2}-667541568n$ (68)
@@
$W^{\mathrm{RW}}_{n,6}(n-10)$ $=$ $\frac{1}{36}(29n^{7}+163n^{6}-28069n^{5}+546697n^{4}-5586400n^{3}+39828436n^{2}-198287652n$ (69)
@@
$W^{\mathrm{RW}}_{n,7}(n-10)$ $=$ $\frac{1}{3}(31n^{6}-332n^{5}-13194n^{4}+342222n^{3}-3412555n^{2}+16788098n$ (70)
@@ Appendix C Polynomials for the number of self-avoiding walks with small slack
$W^{\mathrm{SAW}}_{n,0}(n-10)$ $=$ $\frac{1}{3600}(63n^{10}-6335n^{9}+299060n^{8}-8750150n^{7}+176109699n^{6}-2551326815n^{5}+26957728290n^{4}$ (95)
@@
$W^{\mathrm{SAW}}_{n,1}(n-10)$ $=$ $\frac{1}{72}(35n^{9}-3310n^{8}+144274n^{7}-3810720n^{6}+67350079n^{5}-827584026n^{4}+7081605356n^{3}$ (96)
@@
$W^{\mathrm{SAW}}_{n,2}(n-10)$ $=$ $\frac{1}{36}(195n^{8}-16591n^{7}+636953n^{6}-14439205n^{5}+211951772n^{4}-2069561980n^{3}+13172823368n^{2}$ (97)
@@
$W^{\mathrm{SAW}}_{n,3}(n-10)$ $=$ $\frac{4}{9}(71n^{7}-5124n^{6}+161663n^{5}-2893446n^{4}+31815851n^{3}-216064455n^{2}+846392625n$ (98)
@@
$W^{\mathrm{SAW}}_{n,4}(n-10)$ $=$ $\frac{1}{3}(315n^{6}-17677n^{5}+401955n^{4}-4587843n^{3}+25622178n^{2}-48389612n$ (99)
@@ Appendix D Universality of scaling in internal energy for off-axis directions
$W_{n}^{\mathrm{RW}}(\mathbf{r})=\binom{n}{\frac{n-(r_{x}+r_{y}+r_{z})}{2}}\sum_{k=0}^{\frac{n-(r_{x}+r_{y}+r_{z})}{2}}\binom{\frac{n-(r_{x}+r_{y}+r_{z})}{2}}{k}\binom{\frac{n+r_{x}+r_{y}+r_{z}}{2}}{k+r_{y}+r_{z}}\binom{2k+r_{y}+r_{z}}{k+r_{z}}$. (107)

### v3

@@ I Introduction
[…]
@@ II Models: RW–SAW and SAW–NAW crossovers
[…]
@@
$S(r,\beta\varepsilon)$ $=$ $k_{B}\beta\left[U(r,\beta\varepsilon)-A(r,\beta\varepsilon)\right],$ (5)
@@
[…]
@@ 99 $2069333612$ $32887440$ $204330$
[…]
@@ V Emergence of negative energetic elasticities in lattice polymer chains
[…]
@@ Acknowledgements.
Data Availability
All exact enumeration data used in this study are fully provided within the Supplemental Material. Additional processed data are not publicly available but are available from the authors upon reasonable request.
@@ Appendix A Upper bounds on the number of interacting segment pairs
[…]
$m_{\mathrm{ub}}^{\mathrm{RW}}=\cases{\displaystyle}\left\lfloor\frac{n^{2}}{4}\right\rfloor=\left\lfloor\frac{s^{2}}{4}\right\rfloor&(r=0),\\ \\ \displaystyle\left\lfloor\frac{(n-r+1)^{2}}{4}\right\rfloor=\left\lfloor\frac{(s+1)^{2}}{4}\right\rfloor&(r\geq 1).$ (12)
[…]
@@ Appendix B Polynomials for the number of random walks with small slack
$W^{\mathrm{RW}}_{n,0}(n-8)$ $=$ $\notag\displaystyle\frac{1}{288}(35n^{8}-1620n^{7}+33606n^{6}-407160n^{5}+3151827n^{4}-15991140n^{3}+52024708n^{2}\notag$ (43)
@@
$W^{\mathrm{RW}}_{n,1}(n-8)$ $=$ $\notag\displaystyle\frac{2}{9}(10n^{7}-383n^{6}+6235n^{5}-54875n^{4}+270421n^{3}-654914n^{2}+234858n\notag$ (45)
@@
$W^{\mathrm{RW}}_{n,2}(n-8)$ $=$ $\notag\displaystyle\frac{1}{18}(10n^{7}-182n^{6}+466n^{5}+5671n^{4}+53494n^{3}-1346021n^{2}+7741074n\notag$ (47)
@@
$W^{\mathrm{RW}}_{n,0}(n-10)$ $=$ $\notag\displaystyle\frac{1}{3600}(63n^{10}-4585n^{9}+153060n^{8}-3081950n^{7}+41448599n^{6}-389383445n^{5}+2591570990n^{4}\notag$ (68)
@@
$W^{\mathrm{RW}}_{n,1}(n-10)$ $=$ $\notag\displaystyle\frac{1}{144}(70n^{9}-4395n^{8}+122592n^{7}-1976910n^{6}+20015526n^{5}-127970403n^{4}+476189444n^{3}\notag$ (70)
@@
$W^{\mathrm{RW}}_{n,2}(n-10)$ $=$ $\notag\displaystyle\frac{1}{288}(35n^{9}-970n^{8}-1738n^{7}+356852n^{6}-4644373n^{5}+7548302n^{4}+338801436n^{3}\notag$ (72)
@@
$W^{\mathrm{RW}}_{n,3}(n-10)$ $=$ $\notag\displaystyle\frac{2}{9}(10n^{8}-481n^{7}+10542n^{6}-145552n^{5}+1475385n^{4}-11407483n^{3}+62293463n^{2}\notag$ (74)
@@
$W^{\mathrm{RW}}_{n,4}(n-10)$ $=$ $\notag\displaystyle\frac{1}{18}(5n^{8}+112n^{7}-11966n^{6}+304687n^{5}-4138721n^{4}+34852933n^{3}-185222134n^{2}\notag$ (76)
@@
$W^{\mathrm{RW}}_{n,5}(n-10)$ $=$ $\notag\displaystyle\frac{1}{90}(370n^{7}-7025n^{6}-90077n^{5}+3908575n^{4}-47417645n^{3}+270658210n^{2}-667541568n\notag$ (78)
@@
$W^{\mathrm{RW}}_{n,6}(n-10)$ $=$ $\notag\displaystyle\frac{1}{36}(29n^{7}+163n^{6}-28069n^{5}+546697n^{4}-5586400n^{3}+39828436n^{2}-198287652n\notag$ (80)
@@
$W^{\mathrm{RW}}_{n,7}(n-10)$ $=$ $\notag\displaystyle\frac{1}{3}(31n^{6}-332n^{5}-13194n^{4}+342222n^{3}-3412555n^{2}+16788098n\notag$ (82)
@@ Appendix C Polynomials for the number of self-avoiding walks with small slack
$W^{\mathrm{SAW}}_{n,0}(n-10)$ $=$ $\notag\displaystyle\frac{1}{3600}(63n^{10}-6335n^{9}+299060n^{8}-8750150n^{7}+176109699n^{6}-2551326815n^{5}+26957728290n^{4}\notag$ (108)
@@
$W^{\mathrm{SAW}}_{n,1}(n-10)$ $=$ $\notag\displaystyle\frac{1}{72}(35n^{9}-3310n^{8}+144274n^{7}-3810720n^{6}+67350079n^{5}-827584026n^{4}+7081605356n^{3}\notag$ (110)
@@
$W^{\mathrm{SAW}}_{n,2}(n-10)$ $=$ $\notag\displaystyle\frac{1}{36}(195n^{8}-16591n^{7}+636953n^{6}-14439205n^{5}+211951772n^{4}-2069561980n^{3}+13172823368n^{2}\notag$ (112)
@@
$W^{\mathrm{SAW}}_{n,3}(n-10)$ $=$ $\notag\displaystyle\frac{4}{9}(71n^{7}-5124n^{6}+161663n^{5}-2893446n^{4}+31815851n^{3}-216064455n^{2}+846392625n\notag$ (114)
@@
$W^{\mathrm{SAW}}_{n,4}(n-10)$ $=$ $\notag\displaystyle\frac{1}{3}(315n^{6}-17677n^{5}+401955n^{4}-4587843n^{3}+25622178n^{2}-48389612n\notag$ (116)
@@ Appendix D Universality of scaling in internal energy for off-axis directions
[…]